# PlanetServe: A Decentralized, Scalable, and Privacy-Preserving Overlay for Democratizing Large Language Model Serving

### v4

@@ 5.3 Latency of Confidential Computing (CC)
In Table 1, we show the results of the experiments on an Azure Standard NCC40ads H100 v5 virtual machine (Confidential VM), equipped with a single NVIDIA H100 GPU, 40 vCPUs, and 320 GB RAM, for the CC-on setting. For the CC-off setting, we used the corresponding Standard NC40ads H100 v5 VM, which provides the same hardware resources but does not enable confidential computing features. In both settings, the request rate was fixed at 20 requests per second. The results show that CC adds very small latency.
@@ 5.4 Overlay Forwarding
We evaluate three types of latency of the model node responding to user requests: average generation latency capped at 1000 tokens (Avg), 99th percentile (P99), and Time to First Token (TTFT). The LLMs used in the evaluations are Meta Llama 3 (8B) running on A6000 and DeepSeek-R1 (14B) running on A100. We compare PlanetServe with a centralized baseline of 8 A100 GPUs without KV-cache sharing. Each data point reported here is the average of three independent runs of the entire workload. Latency results are shown in Fig. 14 for PlanetServe and the centralized baseline. PlanetServe consistently demonstrates significant latency improvements across all evaluated workloads, LLM, and GPU setups. Notably, under heavy workload (LooGLE and Mixed), the difference is more evident, indicating the effectiveness of PlanetServe’s overlay forwarding and cache-reuse strategy. The TTFT metric directly impacts perceived user responsiveness, a critical factor in practical deployments. PlanetServe provides significantly shorter latency in TTFT. For both LLMs, PlanetServe reduces TTFT at higher request rates by 40-50%. In addition, for higher request rates, PlanetServe maintains slower latency growth.
To better understand the benefits of different design components, we conduct an ablation study by incrementally enabling the HR-tree and load balancing feature using 8 A100 as 8 model nodes running Llama-3.1-8B. The results are shown in Fig. 17 with the ToolUse workload and a Zipf-1.1 distribution. We find that using the HR-tree almost reduces both the average and P99 latency by more than half, and adding load balancing further reduces latency.
@@ 5.5 Verification throughput
Assuming each model node requires verification 50 times per day and a deployment ratio of up to 100 model nodes per VN, the required throughput is  208 verifications per verification node per hour. We evaluate the verification nodes on two different hardware platforms. On a high-end setup equipped with an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD, a node achieves a throughput of 45.04 verifications per minute. On a standard configuration featuring an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, the verification rate is 20.72 verifications per minute. Both hardware platforms are sufficient to complete the verification tasks.
@@ 6 Discussion
Lesson learned from past projects. We investigate the challenges of the past decentralized overlay networks, such as PlanetLab and P2P file sharing, and discuss how the design of PlanetServe can learn from these projects. 1) Incentive design is important. Performance efficiency alone does not necessarily lead to adoption. Hence, we design a comprehensive incentive infrastructure with verification services. We plan to keep improving this infrastructure. 2) Churn-resilience and security should be considered at the beginning of deployment, which are the main research objectives of this work. 3) Integration with modern computing architecture. Some past P2P networks cannot be integrated with the cloud architecture, which significantly limits their applications. The design of PlanetServe aims at a flexible and scalable architecture that is compatible with cloud and can evolve with new technologies.
More complex models. This paper focuses on text-only tasks. Future research will extend the architecture to support more complex models. A key direction is serving multimodal models including requests with non-textual data. In fact, HR-Tree provides advantages in supporting non-textual data compared to existing KV cache sharing methods, because it records the hash values rather than the specific texts. Another future research direction is to support advanced reasoning and tool-using models. The workloads of these models could be different from the current ones, requiring adaptation of PlanetServe’s current KV cache sharing and load-balancing mechanisms. However, the main infrastructure of PlanetServe can remain the same. New optimization methods can be deployed in it incrementally.
Model partitioning. Due to the limited size of a research paper, we do not consider model partitioning [borzunov2023petal], which can be complementary to PlanetServe. The new combined research objective, i.e., how to achieve decentralized overlay LLM serving with model partitioning, can be future research.
Blockchain integration. Our committee-based BFT mechanism provides a lightweight and efficient solution for serving and verification without incurring the additional overhead of deploying blockchain. However, blockchain offers desirable properties such as transparency, auditability, and simplified global coordination. In particular, smart contracts could automate reputation management in a transparent and tamper-resistant manner. Thus, We will integrate blockchain in the future to enhance transparency and auditability as the system scales to support a larger number of users and model nodes.
@@ 7 Related Works
With the growing importance and popularity of LLM applications, LLM serving engines, software frameworks to run LLM inferences efficiently, have been studied and developed [vllm, PowerInfer, sglang24, Speculative, Preble25, TensorRT, TGI, Fairness2024Sheng, LoongServe2024, DistServe2024, SplitWise2024, CacheGen2024, liu2024droidspeak, InfiniGen]. One key challenge of an LLM serving engine is the management of KV cache. For example, vLLM [vllm] stores KV cache as non-continuous blocks, each with a fixed number of tokens. It enables sharing KV caches for different user requests with similar prompts.
@@ 8 Conclusion
This work introduces PlanetServe, a novel decentralized infrastructure designed for scalable and privacy-preserving LLM serving. We present designs addressing four key challenges: overlay network organization, anonymous communication, overlay forwarding for efficiency, and verification of model serving quality. These issues are fundamental and inherent to any future decentralized LLM-serving system. We implement a prototype of PlanetServe and evaluation results show that PlanetServe achieves low serving latency with minimal overhead for security and trustworthy features. In the future, we plan to deploy PlanetServe at scale and encourage broader community engagement in the research and use of PlanetServe. We believe this work offers a new direction that will attract further research and eventually help democratize future AI innovations and deployment.
@@
A3. The Sentry algrotihm for the chunk length array
@@
Figure 21: Latency of LLM serving with A1000 GPUs
@@
A7. Latency of LLM serving with A1000 GPUs
We evaluate the latency of LLM serving with A1000 GPUs and show the results in Fig. 21. Compared to the A100 results in Fig. 14, PlanetServe shows similar advantages.
@@
An intersection attack [Intersection-CCS13] can be launched to anonymity systems if the target uses pseudonymity and the attacker has access to the user’s history of visits and ISP logs. PlanetServe treats each sequence of prompts as independent. Hence users do not use pseudonymity and the attacker cannot obtain user’s history. Even though the attacker can access ISP logs (very big assumption!), they cannot match the user behavior among the massive amount of Internet users.
@@
Even at inter-continental scales, these latencies are modest compared to LLM inference time (see Fig. 14), suggesting that communication does not dominate end-to-end serving time.

### v5

@@ 5.3 Latency of Confidential Computing (CC)
Table 1 reports results on Azure Standard NCC40ads H100 v5 VMs with CC enabled and the corresponding NC40ads H100 v5 VMs without CC, both equipped with a single NVIDIA H100 GPU, 40 vCPUs, and 320 GB RAM at 20 requests per second. With the request rate fixed at 20 requests per second in both settings, CC introduces minimal latency overhead.
@@ 5.4 Overlay Forwarding
We evaluate three types of latency for the model node responding to user requests: average generation latency capped at 1000 tokens (Avg), 99th percentile (P99), and Time to First Token (TTFT). The LLMs used in the evaluations are Meta Llama 3 (8B) running on A6000 and DeepSeek-R1 (14B) running on A100. We compare PlanetServe with a centralized baseline of 8 A100 GPUs without KV-cache sharing. Each data point reported here is the average of three independent runs of the entire workload. Latency results are shown in Fig. 14 for PlanetServe and the centralized baseline. A detailed comparison against centralized serving with and without KV-cache sharing is provided in Fig. 22 in Appendix A. A8. Latency Comparison with Centralized Baselines. PlanetServe consistently demonstrates significant latency improvements across all evaluated workloads, LLM, and GPU setups. Notably, under heavy workload (LooGLE and Mixed), the difference is more evident, indicating the effectiveness of PlanetServe’s overlay forwarding and cache-reuse strategy. The TTFT metric directly impacts perceived user responsiveness, a critical factor in practical deployments. PlanetServe provides significantly shorter latency in TTFT. For both LLMs, PlanetServe reduces TTFT at higher request rates by 40-50%. In addition, for higher request rates, PlanetServe maintains slower latency growth.
To isolate design benefits, we perform an ablation study by incrementally enabling HR-tree and load balancing on 8 A100 GPUs as model nodes running Llama-3.1-8B. Figure 17 shows results under the ToolUse workload with a Zipf-1.1 distribution. HR-tree reduces both average and P99 latency by over 50%, with load balancing providing further gains.
@@ 5.5 Verification throughput
Assuming each model node requires verification 50 times per day and a deployment ratio of up to 100 model nodes per VN, the required throughput is  208 verifications per VN per hour. We evaluate the verification nodes on two platforms. A high-end GH200 system achieves 45.04 verifications per minute, while a standard A100 system reaches 20.72 per minute. Both configurations meet the required verification throughput.
@@ 6 Discussion
Lesson learned from past projects. We investigate the challenges of past decentralized overlay networks, such as PlanetLab and P2P file sharing, and discuss how the design of PlanetServe can learn from these projects. 1) Incentive design is important. Performance efficiency alone does not necessarily lead to adoption. Hence, we design a comprehensive incentive infrastructure with verification services. 2) Churn resilience and security should be considered at the beginning of deployment, which are the main research objectives of this work. 3) Integration with modern computing architecture. Some past P2P networks cannot integrate with the cloud architecture, which significantly limits their applications. The design of PlanetServe aims at a flexible and scalable architecture compatible with cloud and can evolve with new technologies.
More complex models. This paper focuses on text-only tasks. Future research can extend the architecture to support more complex models. One direction is multimodal support. HR-Tree naturally supports non-textual data by storing hash values rather than raw text, offering advantages over existing KV cache sharing methods. Another is to support advanced reasoning and tool-using models, whose workloads may require adapting PlanetServe’s current KV cache sharing and load-balancing mechanisms. The core PlanetServe infrastructure remains unchanged and can be incrementally optimized.
Model partitioning. Due to space constraints, we do not consider model partitioning [borzunov2023petal], which is complementary to PlanetServe. This combined direction is left for future work.
Blockchain integration. Our committee-based BFT mechanism offers efficient serving and verification without blockchain overhead. Blockchain offers transparency, auditability, and global coordination via smart contracts, and can be integrated to support large-scale deployments.
Practical deployment considerations. We validate PlanetServe’s mechanisms through moderate-scale deployments and large-scale simulations capturing heterogeneous failures. The architecture scales without centralized control, though larger deployments may introduce additional trade-offs.
7 Limitations
Model integrity verification using trusted execution environments. PlanetServe uses confidential computing to protect prompt privacy, which is orthogonal to the model verification mechanism. Hardware-based attestation could strengthen model integrity by verifying that specific model weights are loaded within a trusted execution environment. Prior work such as AttestLLM [zhang2025attestllm] explores this direction for on-device, single-node LLM deployments, but does not address distributed, multi-node, or decentralized serving environments.
Functional alignment verification and model identity. PlanetServe’s verification prioritizes functional alignment, validating whether responses match the expected behavior and quality of the intended model rather than enforcing guarantees on model identity. Thus, it cannot distinguish between different models that produce statistically similar outputs. This limits use cases such as benchmarking that require attributing outputs to a specific implementation. Ensuring model identity without hardware support remains an open challenge.
@@ 8 Related Works
LLM serving engines and software frameworks have been studied and developed [vllm, PowerInfer, sglang24, Speculative, Preble25, TensorRT, TGI, Fairness2024Sheng, LoongServe2024, DistServe2024, SplitWise2024, CacheGen2024, liu2024droidspeak, InfiniGen]. One key challenge of an LLM serving engine is the management of KV cache. For example, vLLM [vllm] stores KV cache as non-continuous blocks, each with a fixed number of tokens. It enables sharing KV caches for different user requests with similar prompts.
@@ 9 Conclusion
This work introduces PlanetServe, a novel decentralized infrastructure for scalable and privacy-preserving LLM serving. We present designs addressing four key challenges: overlay network organization, anonymous communication, overlay forwarding for efficiency, and verification of model serving quality. These issues are fundamental and inherent to future decentralized LLM-serving system. We implement a prototype of PlanetServe and evaluation results show that PlanetServe achieves low serving latency with minimal overhead for security and trustworthy features. We believe this work offers a new direction that will attract further research and eventually help democratize future AI innovations and deployment.
10 Acknowledgement
We thank our shepherd Walter Willinger and the anonymous reviewers for their suggestions and comments. The authors were partially supported by National Science Foundation Grants 2322919, 2420632, 2426031, and 2426940.
@@ 10 Acknowledgement
A3. The Sentry algorithm for the chunk length array
@@
Figure 21: Latency of LLM serving with A100 GPUs
@@
A7. Latency of LLM serving with A100 GPUs
We evaluate the latency of LLM serving with A100 GPUs and show the results in Fig. 21. Compared to the A100 results in Fig. 14, PlanetServe shows similar advantages.
@@
An intersection attack [Intersection-CCS13] can be launched to anonymity systems if the target uses pseudonymity and the attacker has access to the user’s history of visits and ISP logs. PlanetServe treats each sequence of prompts as independent. Hence, users do not use pseudonymity and the attacker cannot obtain user’s history. Even though the attacker can access ISP logs (very big assumption!), they cannot match the user behavior among the massive amounts of Internet users.
@@
Even at inter-continental scales, these latencies are modest compared to LLM inference time (see Fig. 14), suggesting that communication does not dominate end-to-end serving time. Nevertheless, this overhead reflects an inherent privacy–latency trade-off and may still be perceptible in latency-sensitive interactive settings.